Risultati Prova Intermedia di Metodi Matematici C.L. Economia Aziendale del 07/12/05

Cognome Nome Matricola Ex.1 Ex.2 Ex.3 Ex.4 Ex.5 RisultatiRisultati
GAINI ELISA 4170774  0  7  0  0  5 10.2  10
CAMPOLMI DARIO 4154829 10 10 10 10  7 32.4  29
MORETTIALESSIA4168951 10 10  4  0  5 17.2  17
DI GERONIMO MARCO4161997 10  0 10  6  6  22  22
CARRAI SARA 4150473 10  5 10  0  6 20.2  20
MONTANINO ROSA MARIA 4152726  0  5  0  2  4  9.4  9
FARAONIGIORGIA4155384 10 10 10 10  10  36  30
RUTILI MARTA4237150  8  0  0  0  5  7.6  8
MAZZONILEONARDO d.n. 24/10/86  0  2  0  0  0  1.2  1
SPINELLINICO 4165454  6 10  2  2  3  14  14
RAMAZZOTTI LORENZO 3834793 10  3  0  0  5  9.8  10